Electrons are ejected from sodium metal by any light that has a wavelength shorter than 544 nm. What is the kinetic energy of an electron (in J) that is ejected from sodium by light with a wavelength of 325 nm?
Answer #1

wavelength = 544 nm = 544 * 10^-9 m = 5.44 * 10^-7 m

Threshold energy = hc / wavelength = 6.626 * 10^-34 * 3.0 * 10^8 / 5.44 * 10^-7 = 3.654 * 10^-19 J

energy of light = hc / wavelength = 6.626 * 10^-34 * 3.0 * 10^8 / 3.25 * 10^-7 = 6.116 * 10^-19 J

and

kinetic energy of an electron = 6.116 * 10^-19 J - 3.654 * 10^-19 J = 2.46 * 10^-19 J (answer)